# Vaccination in Metropolitan Areas: Should the Capital Be Prioritized in Vaccine Allocation?

### v1

@@ 2 Methods
The formulation of the vaccination model and the associated optimal control problem is composed of three elements: (i) the epidemiological model that combines a compartmental SIR model and commuting patterns to describe the transmission dynamics; (ii) an objective function to be minimized, which contemplates the number of infections and applied vaccines; (iii) a set of constraints on the state and control variables that ensure resource limitations for the production, distribution, and application of the vaccines. For the epidemiological model, we provide bounds for the basic reproduction number.
@@ 2.1 Epidemiological modeling
where $\beta_{i}$ represents city $i$’s infection rate, which is directly affected by the population density and the probability of infection given a contact, and $\gamma^{-1}$ refers to the average infection period.
@@
$\frac{dS_{i}}{dt}(t)&=\mu-\alpha(t)\beta_{i}S_{i}I_{i}-(1-\alpha(t))S_{i}\sum_{j=1}^{K}\beta_{j}p_{ij}I_{j}^{\mathrm{eff}}-\mu S_{i},\\ \frac{dI_{i}}{dt}(t)&=\alpha(t)\beta_{i}S_{i}I_{i}+(1-\alpha(t))S_{i}\sum_{j=1}^{K}\beta_{j}p_{ij}I_{j}^{\mathrm{eff}}-\gamma I_{i}-\mu I_{i},\\ \frac{dR_{i}}{dt}(t)&=\gamma I_{i}-\mu R_{i},\end{aligned}\qquad\qquad\text{for }i=1,\dots,K,$ (1)
subject to the initial conditions $S_{i}(0)=S_{i0},I_{i}(0)=I_{i0}$ and $R_{i}(0)=R_{i0}$, for $i=1,\dots,k$. We use the following vector notations $\boldsymbol{S}(t)=(S_{1}(t),\dots,S_{K}(t))\in\mathbb{R}^{K}$, $\boldsymbol{I}(t)=(I_{1}(t),\dots,I_{K}(t))\in\mathbb{R}^{K}$ and $\boldsymbol{R}(t)=(R_{1}(t),\dots,R_{K}(t))\in\mathbb{R}^{K}$. Figure 1 represents this model graphically.
@@
In this study, we do not consider disease-specific mortality. This decision is based on two observations: firstly, the inclusion of an additional mortality rate due to the disease does not change the disease dynamics. Secondly, our primary focus lies in analyzing the impact of vaccination strategies during the initial months of the epidemic, a period where the relative impact of disease-specific mortality on the outcomes is usually minimal. Our main objective is to control the rate of new infections, which is not influenced by the mortality rate.
@@
is invariant under system (1). The complete statement and its proof are in the Appendix B
Figure 1: Graphical representation of city interactions and the SIR model: The cities are connected composing a network, $p_{ij}$ being the proportion of individuals from city $i$ who work during the day at $j$. At night, cities do not interact, and the epidemic follows a standard SIR dynamic. During the day, some individuals commute, interact with infected individuals in the destination city, and may transport the infection.
@@ 2.2 The basic reproduction number
Following the method develop by Van den Driessche and Watmough [39], we derive the basic reproduction number $\mathcal{R}_{0}$, using the formula $\mathcal{R}_{0}=\rho(FV^{-1})$. Here $F$ and $V$ are matrices associated with the Jacobian of the dynamics at the disease-free equilibrium (DFE). The detailed definitions are included in the Appendix A and the derivation is given in Subsection A.1 Given the problem’s dimension, we could not obtain an explicit expression for $\mathcal{R}_{0}$. However, we obtained in Theorem 2.1, Corollary 2.1 and Theorem 2.2 below, sharp inequalities for $\mathcal{R}_{0}$ in terms of simple expressions involving the parameters.
@@
We get the following result, the proof of which is in the Appendix B.
@@
The basic reproduction number $\mathcal{R}_{0}$ for system (1) satisfies the following inequalities
@@
Figure 2: A version of Figure 1 considering a metropolitan area and the Assumption 2.1.
@@
In a metropolitan area, there is a capital, which is a big city with higher population density and a larger number of inhabitants, along with other connected cities. We assume that individuals residing in the capital city both stay and work there, while people in other cities either commute to the capital or work in their home city. We further enumerate the cities according to their populations in decreasing order of population size. Consequently, matrix $P$ reduces to the form:
@@
When applying the specific structure of Assumption 2.1 to the dynamics described in equation (1), we derive a different inequality for the reproductive number $\mathcal{R}_{0}$ (the proof is in the Appendix B).
Figure 3: Transition matrix of the Rio de Janeiro metropolitan area: a heatmap showing the matrix $P$ in the case of Rio de Janeiro metropolitan area (mobility data taken from [34]).
@@
The basic reproduction number $\mathcal{R}_{0}$ for system (1) under Assumption 2.1, satisfies the following inequalities
@@
where $\xi=\max_{1\leq i\leq K}\alpha\beta_{i}+(1-\alpha)p_{ii}\beta_{i}(1-\delta_{i1})$, where $\delta_{i1}=1$ if, and only if, $i=1$ and $\delta_{i1}=0$ otherwise.
@@
It is evident that higher values of $\gamma$ correspond to faster recovery rates and lower values of $\mathcal{R}_{0}$. As $\alpha$ balances day and night periods, when it approaches $1$, the night period becomes more relevant, and $\mathcal{R}_{0}$ tends to the maximum of $\mathcal{R}_{0}^{i}$, for $i=1,\dots,K$, meaning that infections in the residence city governed the dynamics. The parameters $\beta_{i}$ directly amplify $\mathcal{R}_{0}$. Through numerical experiments, we observed that the highest value among the $\beta_{i}$ parameters is the most significant driver (see e.g. Figure 7). Finally, although the individual population sizes $n_{i}$ do not directly influence eigenvalue behaviors, the ratios $n_{i}/n_{1}$, offer a subdued effect, predominantly guided by $(1-\alpha)\max_{1\leq i\leq K}\beta_{i}$.
@@ 2.3 Vaccination and optimal control problem
$\mathcal{R}_{0}^{\text{vac},i}=\frac{\mu}{\mu+u_{i}}\mathcal{R}_{0}^{i}=\frac{\mu}{\mu+u_{i}}\frac{\beta}{\gamma+\mu}.$
@@
$D(t):=\begin{cases}\displaystyle\sum_{n=0}^{w-1}V_{n}^{\max}+V_{w}^{\max}(t-7w),&t\in[7w,7w+1],\\ \displaystyle\sum_{n=0}^{w}V_{n}^{\max},&t\in[7w+1,7(w+1)],\end{cases}\quad\text{for }w=0,1,\dots,T/7-1,$
where $T$ is a multiple of $7$ for simplicity. The function $D$ is piecewise linear non-decreasing. Note that, to account for the delay in delivering vaccines across cities and healthcare centers, $D$ increases linearly on the first day of each week, after which it remains constant for the rest of the week. This induces the following constraint:
@@
Figure 4: Example of the function of delivered vaccines
@@ 3.1 Simulations for the basic reproduction number
Before we introduce vaccination, we examine the impact of the parameters on the basic reproduction number $\mathcal{R}_{0}$, as described in Section 2 and the bounds for $\mathcal{R}_{0}$ provided by the aforementioned Theorems 2.1 and 2.2. For all the simulations throughout this article, we fix $\gamma$ to $1/7$, which means that the recovery time from infection is $7$ days on average, and $\mu=3.6\cdot 10^{-5}$ to represent approximate the average life expectancy of $75$ years.
@@
Figure 5: Dynamics of the Epidemic in Two Cities: epidemic dynamics in two cities under two different scenarios. The settings are $\alpha=0.64,p_{21}=0.2$ and $n_{1}=10$. The proportion of susceptible individuals is plotted in blue, and the proportion of infectious individuals in orange. The larger city is represented by a solid line, while the smaller city is represented by a dashed line. The basic reproduction number $\mathcal{R}_{0}$, is calculated for each scenario.
In Figure 6 we show the level curves of $\mathcal{R}_{0}$ when $p_{21},$ $n_{1}/n_{2}$ and $\alpha$ vary. For the chosen parameters $\beta,\gamma$ and $\mu$, we have $\mathcal{R}_{0}^{1}>1$ and $\mathcal{R}_{0}^{2}<1,$ which represents a situation for which Corollary 2.1 cannot determine whether $\mathcal{R}_{0}$ is greater or smaller than 1. Looking at Figure 6, we notice that higher values of $\alpha$ increase the value of $\mathcal{R}_{0}$. The impact of $p_{21}$ is nonlinear and complex, but its behavior is similar to a quadratic function in $\mathcal{R}_{0}$. On the other hand, the impact of $p_{21}$ and $\alpha$ is small when the population proportion $n_{1}/n_{2}$ increases. Additionally, see the histogram of Figure C.21 in the Appendix, which illustrates the situation where $\mathcal{R}_{0}^{1}$ is close to 1.
Figure 6: Basic reproduction number as function of $\alpha$ and $p_{21}$ for two cities: contour plot of the values of $\mathcal{R}_{0}$ for different values of $\alpha\in(0,1)$, $p_{21}\in(0,0.9)$ and $n_{1}/n_{2}=(1,2,5,10)$ considering $\beta=(0.4,0.1)$.
@@
Figure 7: Impact of the parameters on the bounds of $\mathcal{R}_{0}$: Using the same setting as in Figure 5, we visualize the impact of $\alpha,p_{21},\beta_{1}$ and $\beta_{2}$ on $\mathcal{R}_{0}$ — depicted in black. In orange, we see the bounds from Theorem 2.2, and in blue, the bounds from Theorem 2.1. The dotted lines are the lower bounds while the dashed lines are the upper bounds.
@@
Our findings, summarized in Table C.2 in the Appendix and illustrated in Figure 8, reveal that, in the capital city, compared to the other cities, the peak size is higher, the peak day occurs earlier, the outbreak [^1] The end of the outbreak is defined when the proportion of infectious individuals is under $10^{-5}$. lasts less and the attack rate is larger. Additionally, higher commuting rates correlate with a shorter epidemic span, as scenarios (II) and (III) exhibit and the capital’s peak day tends to dictate the overall trajectory of the disease spread across the network. Further, the outcomes of scenarios (II) and (III) are similar, which indicates that depending on the structure of the network, we may approximate it using the metropolitan hypothesis in Assumption 2.1.
Figure 8: Simulation results for different structures of the transition matrix: comparison of Peak Size (maximum proportion of infectious individuals), Peak Day (day of the peak size), Duration (time from epidemic onset to the day when the proportion of infectious individuals achieves $10^{-5}$), and Attack Rate (proportion of individuals who contract the disease during $350$ days of the epidemic) across different mobility scenarios. The aggregated data is the weighted average, considering their population size, from cities $2$ to $5$. For this experiment, $\beta=(0.4,0.25,0.2,0.15,0.1)$, $\alpha=0.64$ and the population sizes are $10^{5}\cdot(50,10,10,1,1)$. Lastly, the initial conditions are $I_{1}(0)=I_{2}(0)=I_{3}(0)=10^{-4}$ and no recovered individuals.
@@
Figure 9: Infectious curves for different scenarios of $\beta$: this figure shows the curves of infectious individuals from cities $1$, $3$, and $5$ under four different scenarios of the $\beta$ specification. Here $\alpha=0.64,p_{i1}=0.2$ for $i\geq 2$ and $n_{1}=5n_{2}=5n_{3}=50n_{4}=50n_{5}=5\cdot 10^{6}$. The four scenarios considered are (I) $\beta=(0.2,0.1,0.1,0.1,0.1)$, indicating $\mathcal{R}_{0}^{1}>1$ and $\mathcal{R}_{0}^{i}<1$ for $i\neq 1$; (II) $\beta=(0.15,0.1,0.1,0.1,0.1)$, a situation similar to (I), but with a slower transmissibility rate, even though $\mathcal{R}_{0}>1$; (III) $\beta=(0.1,0.1,0.15,0.1,0.1)$, a situation similar to (II), but with the highest transmissibility rate not in the capital. This scenario results in $\mathcal{R}_{0}<1$; (IV) $\beta=(0.2,0.2,0.2,0.2,0.2)$, where all cities have the same transmissibility rate.
@@ 3.2 Constant vaccination rates
In this scenario, we consider an ongoing epidemic that is subject to a vaccination strategy to control it, starting on a specified day $s$. In the two-city scenario, we set $\beta=(0.5,0.3)$ to ensure that $\mathcal{R}_{0}^{1}>1$ and $\mathcal{R}_{0}^{2}>1$. The vaccination rates are varied within the range $u=(u_{1},u_{2})in\in{[10^{-6},10^{-3}]}^{2}$ to generate Figure 10, which is presented on the $\log$-$\log$ scale. For this setting, the basic reproduction number without vaccination is $\mathcal{R}_{0}\approx 3.48$. Numerically, by a fitting method, we observed that
@@
Figure 10: Basic reproduction number $\mathcal{R}_{0}^{\mathrm{vac}}$ in the vaccination model: this contour plot illustrates the basic reproduction number $\mathcal{R}_{0}^{\mathrm{vac}}$ for the model described in (5) as a function of $u_{1}$ and $u_{2}$ across $K=2$ cities. The settings are: $\beta=(0.5,0.3)$, $\alpha=0.64$, $p_{21}=0.2$ and the ratio $n_{1}/n_{2}=10$, similar to the settings in Figure 6. Both axes $u_{1}$ and $u_{2}$ are in the logarithmic scale.

### v2

# Optimal vaccination strategies on networks and in metropolitan areas

@@ 2 Methods
The formulation of the vaccination model and the associated optimal control problem is composed of three elements: (I) the epidemiological model that combines a compartmental SIR model and commuting patterns to describe the transmission dynamics; (II) an objective function to be minimized, which contemplates the number of infections and applied vaccines; (III) a set of constraints on the state and control variables that ensure resource limitations for the production, distribution, and application of the vaccines.
@@ 2.1 Epidemiological modeling
where $\beta_{i}$ represents city $i$’s infection rate, which is directly affected by the population density and the probability of infection given contact, and $\gamma^{-1}$ refers to the average infection period.
@@
$\frac{dS_{i}}{dt}(t)=\mu-\alpha(t)\beta_{i}S_{i}I_{i}-(1-\alpha(t))S_{i}\sum_{j=1}^{K}\beta_{j}p_{ij}I_{j}^{\mathrm{eff}}-\mu S_{i},$ (1)
$\frac{dI_{i}}{dt}(t)=\alpha(t)\beta_{i}S_{i}I_{i}+(1-\alpha(t))S_{i}\sum_{j=1}^{K}\beta_{j}p_{ij}I_{j}^{\mathrm{eff}}-\gamma I_{i}-\mu I_{i},$
$\frac{dR_{i}}{dt}(t)=\gamma I_{i}-\mu R_{i},$
for $i=1,\dots,K$, subject to the initial conditions $S_{i}(0)=S_{i0},I_{i}(0)=I_{i0}$ and $R_{i}(0)=R_{i0}$, for $i=1,\dots,k$. We use the following vector notations $\boldsymbol{S}(t)=(S_{1}(t),\dots,S_{K}(t))\in\mathbb{R}^{K}$, $\boldsymbol{I}(t)=(I_{1}(t),\dots,I_{K}(t))\in\mathbb{R}^{K}$ and $\boldsymbol{R}(t)=(R_{1}(t),\dots,R_{K}(t))\in\mathbb{R}^{K}$. Figure 1 represents this model graphically.
Figure 1: Graphical representation of city interactions and the SIR model.
The cities are connected composing a network, $p_{ij}$ being the proportion of individuals from city $i$ who work during the day at $j$. At night, cities do not interact, and the epidemic follows a standard SIR dynamic. During the day, some individuals commute, interact with infected individuals in the destination city, and may transport the infection.
@@
In this study, we do not consider disease-specific mortality. This decision is based on two observations: firstly, including an additional mortality rate due to the disease does not change the disease dynamics. Secondly, our primary focus lies in analyzing the impact of vaccination strategies during the initial months of the epidemic, a period where the relative impact of disease-specific mortality on the outcomes is usually minimal. Our main objective is to control the rate of new infections, which is not influenced by the mortality rate.
@@
is invariant under system (LABEL:eq:robot-dance-model). The complete statement and its proof are in Appendix B.
@@ 2.2 The basic reproduction number
Following the method developed by Van den Driessche and Watmough [39], we derive the basic reproduction number using the formula $\mathcal{R}_{0}=\rho(FV^{-1})$. Here $F$ and $V$ are matrices associated with the Jacobian of the dynamics at the disease-free equilibrium (DFE). The detailed definitions are included in the Appendix A and the derivation is given in Subsection A.1. Given the problem’s dimension, we could not obtain an explicit expression for $\mathcal{R}_{0}$. However, we obtained in Theorem 2.1, Corollary 2.1 and Theorem 2.2 below, sharp inequalities for $\mathcal{R}_{0}$ in terms of simple expressions involving the parameters.
@@
We get the following result, the proof of which is in Appendix B .
@@
The basic reproduction number $\mathcal{R}_{0}$ for system (LABEL:eq:robot-dance-model) satisfies the following inequalities
@@
Figure 2: Version of Figure 1 considering a metropolitan area and Assumption 2.1.
@@
In a metropolitan area, there is a capital, which is a big city with a higher population density and a larger number of inhabitants, along with other connected cities. We assume that individuals residing in the capital city both stay and work there, while people in other cities either commute to the capital or work in their home city. We further enumerate the cities according to their populations in decreasing order of population size. Consequently, matrix $P$ reduces to the form:
@@
Figure 3: Transition matrix of the Rio de Janeiro metropolitan area.
A heatmap showing the matrix $P$ in the case of the Rio de Janeiro metropolitan area (mobility data taken from [34]).
When applying the specific structure of Assumption 2.1 to the dynamics described in equation (LABEL:eq:robot-dance-model), we derive a different inequality for the reproductive number $\mathcal{R}_{0}$ (the proof is in Appendix B).
@@
The basic reproduction number $\mathcal{R}_{0}$ for system (LABEL:eq:robot-dance-model) under Assumption 2.1, satisfies the following inequalities
@@
where $\xi=\max_{1\leq i\leq K}\alpha\beta_{i}+(1-\alpha)p_{ii}\beta_{i}(1-\delta_{i1})$, where $\delta_{ij}$ is the Kronecker function.
@@
Higher values of $\gamma$ correspond to faster recovery rates and lower values of $\mathcal{R}_{0}$. As $\alpha$ balances day and night periods, when it approaches $1$, the night period becomes more relevant, and $\mathcal{R}_{0}$ tends to the maximum of $\mathcal{R}_{0}^{i}$, for $i=1,\dots,K$, meaning that infections in the residence city governed the dynamics. The parameters $\beta_{i}$ directly amplify $\mathcal{R}_{0}$. Through numerical experiments, we observed that the highest value among the $\beta_{i}$ parameters is the most significant driver (see e.g. Figure 7). Finally, although the individual population sizes $n_{i}$ do not directly influence eigenvalue behaviors, the ratios $n_{i}/n_{1}$ offer a subdued effect, predominantly guided by $(1-\alpha)\max_{1\leq i\leq K}\beta_{i}$.
@@ 2.3 Vaccination and optimal control problem
$\mathcal{R}_{0}^{\text{vac},i}=\frac{\mu}{\mu+u_{i}}\mathcal{R}_{0}^{i}=\frac{\mu}{\mu+u_{i}}\frac{\beta_{i}}{\gamma+\mu}.$
@@
$D(t):=\begin{cases}\displaystyle\sum_{n=0}^{w-1}V_{n}^{\max}+V_{w}^{\max}(t-7w),&t\in[7w,7w+1],\\ \displaystyle\sum_{n=0}^{w}V_{n}^{\max},&t\in[7w+1,7(w+1)],\end{cases}$
for $w=0,1,\dots,T/7-1$ and $T$ a multiple of $7$ for simplicity. The function $D$ is piecewise linear non-decreasing. Note that, to account for the delay in delivering vaccines across cities and healthcare centers, $D$ increases linearly on the first day of each week, after which it remains constant for the rest of the week. This induces the following constraint:
@@
Figure 4: Example of the function of delivered vaccines.
@@ 3.1 Simulations for the basic reproduction number
Before we introduce vaccination, we examine the impact of the parameters on the basic reproduction number $\mathcal{R}_{0}$, as described in Methods and the bounds for $\mathcal{R}_{0}$ provided by the aforementioned Theorems 2.1 and 2.2. For all the simulations throughout this article, we fix $\gamma$ to $1/7$, which means that the recovery time from infection is on average $7$ days, and $\mu=3.6\cdot 10^{-5}$ to represent approximately the average life expectancy of $75$ years.
@@
Figure 5: Dynamics of the Epidemic in Two Cities.
Epidemic dynamics in two cities under two different scenarios. The settings are $\alpha=0.64,p_{21}=0.2$ and $n_{1}=10$. The proportion of susceptible individuals is plotted in blue, and the proportion of infectious individuals in orange. The larger city is represented by a solid line, while the smaller city is represented by a dashed line.
In Figure 6 we show the level curves of $\mathcal{R}_{0}$ when $p_{21},$ $n_{1}/n_{2}$ and $\alpha$ vary. For the chosen parameters $\beta,\gamma$ and $\mu$, we have $\mathcal{R}_{0}^{1}>1$ and $\mathcal{R}_{0}^{2}<1,$ which represents a situation for which Corollary 2.1 cannot determine whether $\mathcal{R}_{0}$ is greater or smaller than 1. Looking at Figure 6, we notice that higher values of $\alpha$ increase the value of $\mathcal{R}_{0}$. The impact of $p_{21}$ is nonlinear and complex, but its behavior is similar to a quadratic function in $\mathcal{R}_{0}$. On the other hand, the impact of $p_{21}$ and $\alpha$ is small when the population proportion $n_{1}/n_{2}$ increases. Additionally, see the histogram of Figure C.21, which illustrates the situation where $\mathcal{R}_{0}^{1}$ is close to 1.
Figure 6: Basic reproduction number as function of $\alpha$ and $p_{21}$ for two cities.
Contour plot of the values of $\mathcal{R}_{0}$ for different values of $\alpha\in(0,1)$, $p_{21}\in(0,0.9)$ and $n_{1}/n_{2}=(1,2,5,10)$ considering $\beta=(0.4,0.1)$.
@@
Figure 7: Impact of the parameters on the bounds of $\mathcal{R}_{0}$.
Using the same setting as in Figure 5, we visualize the impact of $\alpha,p_{21},\beta_{1}$ and $\beta_{2}$ on $\mathcal{R}_{0}$ (depicted in black). In orange, we see the bounds from Theorem 2.2, and in blue, the bounds from Theorem 2.1. The dotted lines are the lower bounds while the dashed lines are the upper bounds.
@@
Our findings, summarized in Table C.2 and illustrated in Figure 8, reveal that in the capital city – compared to the other cities – the peak size is higher, the peak day occurs earlier, the outbreak (period between $t=0$ and the time at which the proportion of infectious individuals falls below $10^{-5}$) lasts less and the attack rate is larger. Additionally, higher commuting rates correlate with a shorter epidemic span, as scenarios (II) and (III) exhibit and the capital’s peak day tends to dictate the overall trajectory of the disease spread across the network. Further, the outcomes of scenarios (II) and (III) are similar, which indicates that depending on the structure of the network, we may approximate it using the metropolitan hypothesis in Assumption 2.1.
Figure 8: Simulation results for different structures of the transition matrix.
Comparison of Peak Size (maximum proportion of infectious individuals), Peak Day (day of the peak size), Duration (time from epidemic onset to the day when the proportion of infectious individuals achieves $10^{-5}$), and Attack Rate (proportion of individuals who contract the disease during $350$ days of the epidemic) across different mobility scenarios. The aggregated data is the weighted average, considering their population size, from cities $2$ to $5$. For this experiment, $\beta=(0.4,0.25,0.2,0.15,0.1)$, $\alpha=0.64$, and the population sizes are $10^{5}\cdot(50,10,10,1,1)$. Lastly, the initial conditions are $I_{1}(0)=I_{2}(0)=I_{3}(0)=10^{-4}$ and no recovered individuals.
@@
Figure 9: Infectious curves for different scenarios of $\beta$.
Curves of infectious individuals from cities $1$, $3$, and $5$ under four scenarios of the $\beta$ specification. Here $\alpha=0.64,p_{i1}=0.2$ for $i\geq 2$ and $n_{1}=5n_{2}=5n_{3}=50n_{4}=50n_{5}$ $=5\cdot 10^{6}$. The four scenarios considered are (I) $\beta=(0.2,0.1,0.1,0.1,0.1)$, indicating $\mathcal{R}_{0}^{1}>1$ and $\mathcal{R}_{0}^{i}<1$ for $i\neq 1$; (II) $\beta=(0.15,0.1,0.1,0.1,0.1)$, a situation similar to (I), but with a slower transmissibility rate, even though $\mathcal{R}_{0}>1$; (III) $\beta=(0.1,0.1,0.15,0.1,0.1)$, a situation similar to (II), but with the highest transmissibility rate not in the capital. This scenario results in $\mathcal{R}_{0}<1$; (IV) $\beta=(0.2,0.2,0.2,0.2,0.2)$, where all cities have the same transmissibility rate.
@@ 3.2 Constant vaccination rates
In this scenario, we consider an ongoing epidemic that is subject to a vaccination strategy to control it, starting on a specified day $s$. In the two-city scenario, we set $\beta=(0.5,0.3)$ to ensure that $\mathcal{R}_{0}^{1}>1$ and $\mathcal{R}_{0}^{2}>1$. The vaccination rates are varied within the range $u=(u_{1},u_{2})\in{[10^{-6},10^{-3}]}^{2}$ to generate Figure 10, which is presented on the $\log$-$\log$ scale. For this setting, the basic reproduction number without vaccination is $\mathcal{R}_{0}\approx 3.48$. Numerically, by a fitting method, we observe that
@@
Figure 10: Basic reproduction number $\mathcal{R}_{0}^{\mathrm{vac}}$ in the vaccination model.
The basic reproduction number $\mathcal{R}_{0}^{\mathrm{vac}}$ for the model described in (5) as a function of $u_{1}$ and $u_{2}$ across $K=2$ cities. The settings are: $\beta=(0.5,0.3)$, $\alpha=0.64$, $p_{21}=0.2$ and the ratio $n_{1}/n_{2}=10$. Both axes $u_{1}$ and $u_{2}$ are in the logarithmic scale.